OpenStudy (shrutipande9):

find the polar form of i^7+i^10

hartnn (hartnn):

first simplify i^7 and i^10 ?

OpenStudy (shrutipande9):

that would be z=-1-i...rite?

hartnn (hartnn):

yes!

hartnn (hartnn):

know how to get polar form ?

OpenStudy (shrutipande9):

i m just confused at d part where u hv 2 find thita

hartnn (hartnn):

theta = tan inverse y/x

hartnn (hartnn):

theta = tan inverse (-1/-1)

OpenStudy (shrutipande9):

z=r(\[\cos \theta + i\])

OpenStudy (shrutipande9):

ooppss

hartnn (hartnn):

z= r cos theta+ i r sin theta

OpenStudy (shrutipande9):

i mean z=r(cos theta+i sin theta)

hartnn (hartnn):

r =... ?

OpenStudy (shrutipande9):

so i get tan inverse 1 so i m gettin theta as pi/4 but d ans says its 3pi/4

OpenStudy (shrutipande9):

r i got as root 2

hartnn (hartnn):

yes, and i'll tell u the reason you tell me your r is positive or negative ?

OpenStudy (shrutipande9):

positive

hartnn (hartnn):

-1-i = -(1+i)

hartnn (hartnn):

so, - root 2 angle 45

OpenStudy (shrutipande9):

yep

hartnn (hartnn):

but r cannot be negative so, +root 2 angle (45-180)

OpenStudy (shrutipande9):

hold on...45-180 kyu kiya?

OpenStudy (shrutipande9):

it cannot b negative so? -180?

hartnn (hartnn):